# Muffled Murmurs: Environmental effects in the LISA stochastic signal from stellar-mass black hole binaries

### v1

@@ III.3 Phenomenological parametric models of stochastic signals containing environmental effects
The stochastic signals described in Secs. III.1 and III.2 are computationally expensive to evaluate for data analysis purposes. In order to analyze the effects of dynamical friction and accretion, we develop “ready-to-use” phenomenological models. As a first step, we construct a “Rational Power-Law” model $\Omega_{\rm RPL}$ given by
@@
where $A_{\rm vac}$ is the vacuum amplitude, $\gamma$ the vacuum spectral index, $\alpha$ is a phenomenological coefficient, $\{\beta,\kappa\}$ are spectral indices parameterizing each environmental effect, and $A_{m}$ controls the fractional change in the SGWB amplitude due to the environment. We obtain $A_{\rm vac}$ and $A_{m}$ using asymptotic matching at low and high frequencies. Specifically, we expand both $\Omega_{\rm RPL}$ and $\Omega_{\rm GW}$ to leading order for $f_{s}\ll f_{\rm turn}$ and for $f_{s}\gg f_{\rm turn}$ and solve for $A_{\rm vac}$ and $A_{m}$. We provide details of the asymptotic matching in Appendix C, while values for $A_{\rm vac}$ and $A_{m}$ are listed in Table 1.
@@
[…]
@@
Here $A_{g}(\alpha)$, $f_{\mathrm{peak}}(\alpha)$ and $\sigma$ control the amplitude, central log-frequency, and width of the correction to $\Omega_{\rm RPL}$. In Table 1, we show the mapping between model parameters and the specific cases of dynamical friction and gas accretion. We discuss the accuracy of our phenomenological models with respect to the “accurate” stochastic signals from Secs. III.1 and III.2 in Appendix C.
@@ IV Detection and parameter estimation of environmental effects
In this section, we first review how stochastic signals are analyzed in LISA data, then show results for how environmental effects (dynamical friction or accretion) can be measured using the phenomenological models developed in Sec. III.3.
@@ IV.1 Detecting stochastic signals in LISA
We model LISA data as linear, time-delayed combinations of six single-link measurements, known as (TDI) variables [70]. TDIs are employed to suppress laser frequency noise, with different combinations applied depending on satellite orbits assumptions. [71, 72, 73, 70, 74]. In this work, we use uncorrelated AET combinations suitable for a static, equal-arm of the LISA constellation [74].
Moreover, in our data model, we do not account for additional features beyond a stationary Gaussian spectrum such as anisotropy, non-Gaussianity, and non-stationarity. While these characteristics may help distinguish overlapping backgrounds they also significantly increase the complexity of the analysis [75, 76, 77, 78, 79].
@@
×$\times\left[1+\tanh{((f_{\mathrm{kn}}-f)/f_{2})}\right].$ (24)
@@
We simulate data for $S_{\mathrm{GW,sBBH}}$ using $\Omega_{\rm GW,sBBH}$ models introduced in Secs. II and III. Instead, for parameter estimation, we use the RPLP models introduced in Eqs. 17 and 18, parameterized by $\theta_{\rm sBBH}=\{A_{\rm vac},\gamma,\alpha,\beta,\kappa,A_{m}\}$. Thus, in the most general case, the inference parameter space is $\theta=\theta_{n}\cup\theta_{\rm Gal}\cup\theta_{\rm sBBH}$.
@@ IV.2 Measuring environmental effects
[…]
@@
[…]
with $\Pi(\theta)$ being the prior over model parameters $\theta$. We consider $\log_{10}\mathcal{B}_{\rm vac}^{\rm non-vac}>1$ to be strong evidence in favor of the non-vacuum hypothesis. To perform parameter estimation we simulate data and infer upon them using the publicly available codebase Bahamas [87]. To compute the evidence, inference is performed using the nested sampling algorithm [88] as implemented in nessai [89].
@@ IV.2.1 Dynamical friction
We first consider the effects of dynamical friction by generating injected SGWB data (using methods of Sec. III.1) with $\rho^{\rm inj}\in\{10^{-7},10^{-8},10^{-9},10^{-10},10^{-11}\}\mathrm{g\,cm^{-3}}$. For the RPLP model parameters $\theta_{\rm sBBH}$, we are mainly interested in recovering the vacuum amplitude $A_{\rm vac}$ and the environmental parameter $\alpha$, which in this case is just the disk density $\rho$. Since $A_{\rm vac}$ and $\gamma$, are strongly correlated, we fix $\gamma=2/3$ in our inference. Likewise, we also set $\{A_{m},\beta,\kappa\}$ to the values listed in Table 1, due to the strong correlation with $\alpha$. These parameters are therefore not sampled in our parameter estimation experiments.
@@
Figure 3: Marginalized posteriors for the dynamical friction model across various matter density regimes. Solid (dashed) contours and histograms correspond to analyses with (without) the inclusion of the Galactic foreground. Two-dimensional contours correspond to $90\%$ credible regions. Dash-dot represent the true value for $\rho$ and the single asymptotic value for $A_{\rm vac}$ (as listed in Table 1).
In Fig. 3, we show the one-dimensional and two-dimensional marginalized posteriors for $A_{\rm vac}$ and $\rho$, inferred with fixed (dashed curves) or free (solid curves) $\theta_{\rm Gal}$. Overall, we find that the recovered posteriors are well constrained relative to the prior, and the effect of dynamical friction is indeed measurable.
For known Galactic foreground parameters, we observe a trend in the posteriors for increasing $\rho^{\rm inj}$. When $\rho^{\rm inj}\gtrsim 10^{-9}\mathrm{g}\,\mathrm{cm}^{-3}$, the recovered posteriors of $\rho$ do not overlap, implying that we can better resolve between different orders of magnitude of the disk density. We quantify this further by computing the statistical uncertainty $\delta\rho/\rho$ using Eq. 25. As listed in Table 2, with increasing $\rho^{\rm inj}$, we find that $\delta\rho$ decreases. Specifically, we find that $\delta\rho\lesssim\rho^{\rm inj}$ when $\rho^{\rm inj}\gtrsim 10^{-9}\mathrm{g}\,\mathrm{cm}^{-3}$, with known Galactic foreground parameters. In all cases, we can infer the vacuum amplitude $A_{\rm vac}$ to $\mathcal{O}(1)\%$ precision. The fact that we can strongly measure the vacuum amplitude is consistent with the results of [30], which also investigated the detectability of SGWB produced by sBBHs in vacuum. While the one-dimensional marginalized posterior of $\rho$ shrinks with increasing $\rho^{\rm inj}$, the one-dimensional marginalized posterior of $A_{\rm vac}$ widens instead due to the correlation between the two parameters.
@@
We now discuss the impact of the Galactic foreground parameters on the marginalized posteriors shown in Fig. 3. As expected, the simultaneous inference on Galactic foreground parameters results in wider marginal posteriors. For $\rho^{\rm inj}\lesssim 10^{-9}\mathrm{g}\,\mathrm{cm}^{-3}$, however, fixing the Galactic foreground parameters $\theta_{\rm Gal}$ causes an underestimation of $A_{\rm vac}$ and $\rho$ compared to inferring on them. At higher densities, the influence of Galactic foreground becomes negligible. This is related to how the SNR is affected by the Galactic foreground for different densities.
@@
In Fig. 4, we show the cumulative SNR with and without including the Galactic foreground parameters, for different values of $\rho^{\rm inj}$. We compute the SNR of the posterior predictive (see App. C for details). We find that as $\rho^{\rm inj}$ increases, the total SNR decreases, consistent with the behavior of the SGWB shown in Fig. 1. Further, including the Galactic foreground lowers the SNR, which along with the additional five Galactic foreground parameters, contributes to wider posteriors. From Fig. 4, we also observe that with increasing $\rho^{\rm inj}$, the difference in SNR with and without Galactic foreground decreases. This is because as $\rho^{\rm inj}$ increases, the turning point shifts to frequencies above $\sim 5\times 10^{-3}$z, where the Galactic foreground is suppressed. Consequently, the posteriors of $\rho$ with and without Galactic foreground overlap better as $\rho^{\rm inj}$ is increased, with virtually no difference in the case of $\rho^{\rm inj}\gtrsim 10^{-8}\mathrm{g}\,\mathrm{cm}^{-3}$. Additionally, the recovered SNR (with and without Galactic foreground) typically accumulates above $50$ only at $f\gtrsim(\mathrm{few})\times 10^{-3}$z. Thus, the measurements of signal parameters are typically informed by this specific frequency regime.
For each injection, we performed a separate set of runs with a vacuum SGWB model by setting $\alpha=\rho=0$ in $\Omega_{\rm RPLP}$. We compute the Bayes factor $\mathcal{B}^{\rm non-vac}_{\rm vac}$ in favor of the non-vacuum model ($\rho\neq 0$), given by Eq. 27, which we list in Table 2. For all injections, we find that $\mathcal{B}^{\rm non-vac}_{\rm vac}>10$, even when including the Galactic foreground. Thus, despite the marginalized posteriors for $A_{\rm vac}$ and $\rho$ displaying biases, the RPLP model can detect dynamical friction effects by strongly disfavoring the vacuum model. We also checked that the posterior predictive from the RPLP model can accuratetely describe the injected signal within statistical errors, consistent with our Bayes factor results. We discuss further in Appendix C and also comment further on the biases observed in Fig. 3.
@@
If we detect a SGWB consistent with vacuum, our RPLP model can also be used to place upper bounds on the environmental parameters. In Fig. 5, we show the constraints on $\rho$, with and without including the Galactic foreground parameters in our model. The constraints, as given by the $90\%$ one-sided credible interval, are informative, because the prior extends to $\rho=10^{-6}\mathrm{g\,cm^{-3}}$. When excluding the Galactic foreground, we find an upper bound of $\rho<8.51\times 10^{-11}\mathrm{g}\,\mathrm{cm}^{-3}$. Owing to the increase in dimensionality when including Galactic foreground, the constraint weakens slightly to $\rho<7.58\times 10^{-10}\mathrm{g}\,\mathrm{cm}^{-3}$.
Figure 6: Marginalized posterior probability for the vacuum model across different matter density regimes. Solid and dashed histograms represent the inference results with and without the inclusion of the Galactic foreground, respectively. Dash-dotted lines indicate the asymptotic value of $A_{\rm vac}$ (from Table 1) with $\gamma=2/3$. Two-dimensional contours correspond to $90\%$ credible regions.
To explore systematic biases on the vacuum SGWB induced by neglecting environmental effects, we inject an SGWB containing dynamical friction effects, and analyze it with the vacuum model, i.e. $\Omega_{\rm RPL}$ with $\alpha=0$. In Fig. 6, for the injected values of $\rho^{\rm inj}\in\{10^{-7},10^{-8},10^{-9},10^{-10},10^{-11}\}\mathrm{g}\,\mathrm{cm}^{-3}$, we show the recovered two- and one-dimensional marginalized posteriors for the vacuum amplitude $A_{\rm vac}$ and the vacuum spectral index $\gamma$. When assessing the systematic bias in $A_{\rm vac}$ and $\gamma$, we compare the maximum posterior point to the asymptotic vacuum values (vertical dash-dotted lines) listed in Table 1. As expected, with increasing $\rho^{\rm inj}$, the systematic biases increase in significance. In more details, $\gamma$ is biased to values larger than the asymptotic value of $\gamma=2/3$, because the SGWB containing dynamical friction effects has a steeper slope, with an asymptotic value of 13/3. Due to the expected positive correlation with $A_{\rm vac}$ (given the functional form of the SGWB), we find that $A_{\rm vac}$ is biased to larger values. When inferring simultaneously on Galactic foreground parameters, due to the reduced constraining power on dynamical friction show in Fig. 3, the biases on $A_{\rm vac}$ and $\gamma$ are milder, as expected.
@@ IV.2.2 Gas accretion
To characterize gas accretion measurability, we inject SGWB data (using models in Sec. III.2) with $f_{\rm Edd}^{\rm inj}\in\{0.01,0.10,1.00,10\}$. We use a log-uniform prior given by $\log_{10}f_{\rm Edd}\sim\mathcal{U}(-3,2)$, where the upper bound reflects astrophysical expectations of how large the Eddington ratio can be. We have checked that our results are robust to the exact choice of the lower bound, which we cannot set exactly to zero due to the log-uniform prior. We carry out the Bayesian inference just like we did for dynamical friction by fixing $\{\gamma,A_{m},\beta,\kappa\}$.
In Fig. 7, we show the two-dimensional and one-dimensional marginalized posteriors of $A_{\rm vac}$ and $f_{\rm Edd}$. We accurately measure the vacuum amplitude across all injections, thus finding that marginalizing over gas accretion does not impact the measurability of the vacuum stellar SGWB. However, we do not obtain informative posteriors of Eddington ratios for any of the injected values, even with $f^{\rm inj}_{\rm Edd}=10$ marginally recovering the prior. The lack of constraining power on $f_{\rm Edd}$ is due to the turning point of the SGWB occurring at frequencies lower that the LISA sensitive band, as we anticipated qualitatively with BPLSs in Sec. III.2. Given our findings, we do not perform any additional Bayesian analysis for gas accretion effects.

### v2

@@ III.3 Phenomenological parametric models of stochastic signals containing environmental effects
The stochastic signals described in Secs.˜III.1 and III.2 are computationally expensive to evaluate for data analysis purposes. In order to analyze the effects of dynamical friction and accretion, we develop “ready-to-use” phenomenological models. As a first step, we construct a “Rational Power-Law” model $\Omega_{\rm RPL}$ given by
@@
where $A_{\rm vac}$ is the vacuum amplitude, $\gamma$ the vacuum spectral index, $\alpha$ is a phenomenological coefficient, $\{\beta,\kappa\}$ are spectral indices parameterizing each environmental effect, and $A_{m}$ controls the fractional change in the SGWB amplitude due to the environment. We obtain $A_{\rm vac}$ and $A_{m}$ using asymptotic matching at low and high frequencies. Specifically, we expand both $\Omega_{\rm RPL}$ and $\Omega_{\rm GW}$ to leading order for $f_{s}\ll f_{\rm turn}$ and for $f_{s}\gg f_{\rm turn}$ and solve for $A_{\rm vac}$ and $A_{m}$. We provide details of the asymptotic matching in Appendix C, while values for $A_{\rm vac}$ and $A_{m}$ are listed in Table˜1.
@@
$\begin{aligned} \Omega_{\mathrm{RPLP}}=\dfrac{\Omega_{\rm RPL}}{1+\mathcal{G}(f,\alpha)},\end{aligned}$ (18)
@@
Here $A_{g}(\alpha)$, $f_{\mathrm{peak}}(\alpha)$ and $\sigma$ control the amplitude, central log-frequency, and width of the correction to $\Omega_{\rm RPL}$. In Table˜1, we show the mapping between model parameters and the specific cases of dynamical friction and gas accretion. We discuss the accuracy of our phenomenological models with respect to the “accurate” stochastic signals from Secs.˜III.1 and III.2 in Appendix C.
@@ IV Detection and parameter estimation of environmental effects
In this section, we first review how stochastic signals are analyzed in LISA data, then show results for how environmental effects (dynamical friction or accretion) can be measured using the phenomenological models developed in Sec.˜III.3.
@@ IV.1 Detecting stochastic signals in LISA
We model LISA data as linear, time-delayed combinations of six single-link measurements, known as time-delay interferometry (TDI) variables [77]. TDI variables are employed to suppress laser frequency noise, with different combinations applied depending on satellite orbits assumptions. [78, 79, 80, 77, 81]. In this work, we use the uncorrelated TDI variables A, E, and T, suitable for a static, equal-arm of the LISA constellation [81].
Moreover, in our data model, we do not account for additional features beyond a stationary Gaussian spectrum such as anisotropy, non-Gaussianity, and non-stationarity. While these characteristics may help distinguish overlapping backgrounds, they also significantly increase the complexity of the analysis [82, 83, 84, 85, 86].
@@
$\times\left[1+\tanh{((f_{\mathrm{kn}}-f)/f_{2})}\right].$ (24)
@@
We simulate data for $S_{\mathrm{GW,sBBH}}$ using $\Omega_{\rm GW,sBBH}$ models introduced in Secs.˜II and III. Instead, for parameter estimation, we use the RPLP models introduced in Eqs.˜17 and 18, parameterized by $\theta_{\rm sBBH}=\{A_{\rm vac},\gamma,\alpha,\beta,\kappa,A_{m}\}$. Thus, in the most general case, the inference parameter space is $\theta=\theta_{n}\cup\theta_{\rm Gal}\cup\theta_{\rm sBBH}$.
@@ IV.2 Measuring environmental effects
$\mathrm{SNR}=\sqrt{T\sum_{k}4\int_{0}^{\infty}df\left(\frac{S_{\rm k,GW}(f)}{S_{\rm k,n}(f)}\right)^{2}},$ (26)
@@
$\mathcal{Z}(\hat{\bm{S}})=\displaystyle\int d\theta\mathcal{L}(\hat{\bm{S}}|\theta)\Pi(\theta),$ (28)
with $\Pi(\theta)$ being the prior over model parameters $\theta$. We consider $\log_{10}\mathcal{B}_{\rm vac}^{\rm non-vac}>1$ to be strong evidence in favor of the non-vacuum hypothesis. To perform parameter estimation we simulate data and infer upon them using the publicly available codebase Bahamas [94, 95]. To compute the evidence, inference is performed using the nested sampling algorithm [96] as implemented in nessai [97].
@@ IV.2.1 Dynamical friction
We first consider the effects of dynamical friction by generating injected SGWB data (using methods of Sec.˜III.1) with $\rho^{\rm inj}\in\{10^{-7},10^{-8},10^{-9},10^{-10},10^{-11}\}\mathrm{g\,cm^{-3}}$. For the RPLP model parameters $\theta_{\rm sBBH}$, we are mainly interested in recovering the vacuum amplitude $A_{\rm vac}$ and the environmental parameter $\alpha$, which in this case is just the disk density $\rho$. Since $A_{\rm vac}$ and $\gamma$, are strongly correlated, we fix $\gamma=2/3$ in our inference. Likewise, we also set $\{A_{m},\beta,\kappa\}$ to the values listed in Table˜1, due to the strong correlation with $\alpha$. These parameters are therefore not sampled in our parameter estimation experiments.
@@
Figure 3: Marginalized posteriors for the dynamical friction model across various matter density regimes. Solid (dashed) contours and histograms correspond to analyses with (without) the inclusion of the Galactic foreground. Two-dimensional contours correspond to $90\%$ credible regions. Dash-dot represent the true value for $\rho$ and the single asymptotic value for $A_{\rm vac}$ (as listed in Table˜1).
In Fig.˜3, we show the one-dimensional and two-dimensional marginalized posteriors for $A_{\rm vac}$ and $\rho$, inferred with fixed (dashed curves) or free (solid curves) $\theta_{\rm Gal}$. Overall, we find that the recovered posteriors are well constrained relative to the prior, and the effect of dynamical friction is indeed measurable.
For known Galactic foreground parameters, we observe a trend in the posteriors for increasing $\rho^{\rm inj}$. When $\rho^{\rm inj}\gtrsim 10^{-9}\mathrm{g}\,\mathrm{cm}^{-3}$, the recovered posteriors of $\rho$ do not overlap, implying that we can better resolve between different orders of magnitude of the disk density. We quantify this further by computing the statistical uncertainty $\delta\rho/\rho$ using Eq.˜25. As listed in Table˜2, with increasing $\rho^{\rm inj}$, we find that $\delta\rho$ decreases. Specifically, we find that $\delta\rho\lesssim\rho^{\rm inj}$ when $\rho^{\rm inj}\gtrsim 10^{-9}\mathrm{g}\,\mathrm{cm}^{-3}$, with known Galactic foreground parameters. In all cases, we can infer the vacuum amplitude $A_{\rm vac}$ to $\mathcal{O}(1)\%$ precision. The fact that we can strongly measure the vacuum amplitude is consistent with the results of [31], which also investigated the detectability of SGWB produced by sBBHs in vacuum. While the one-dimensional marginalized posterior of $\rho$ shrinks with increasing $\rho^{\rm inj}$, the one-dimensional marginalized posterior of $A_{\rm vac}$ widens instead due to the correlation between the two parameters. We note that the one-dimensional marginalized posteriors for $\rho$ are biased for increasing values of $\rho^{\rm inj}$. In Appendix C, we show that this is due to mismatch between the RPLP model and the true signal caused by the approximate nature of the asymptotic matching used. However, we also show that despite the biases in the marginalized posteriors, the posterior predictive is consistent with the injected signal.
@@
We now discuss the impact of the Galactic foreground parameters on the marginalized posteriors shown in Fig.˜3. As expected, the simultaneous inference on Galactic foreground parameters results in wider marginal posteriors. For $\rho^{\rm inj}\lesssim 10^{-9}\mathrm{g}\,\mathrm{cm}^{-3}$, however, fixing the Galactic foreground parameters $\theta_{\rm Gal}$ causes an underestimation of $A_{\rm vac}$ and $\rho$ compared to inferring on them. At higher densities, the influence of Galactic foreground becomes negligible. This is related to how the SNR is affected by the Galactic foreground for different densities.
@@
In Fig.˜4, we show the cumulative SNR with and without including the Galactic foreground parameters, for different values of $\rho^{\rm inj}$. We compute the SNR of the posterior predictive (see App.˜C for details). We find that as $\rho^{\rm inj}$ increases, the total SNR decreases, consistent with the behavior of the SGWB shown in Fig.˜1. Further, including the Galactic foreground lowers the SNR, which along with the additional five Galactic foreground parameters, contributes to wider posteriors. From Fig.˜4, we also observe that with increasing $\rho^{\rm inj}$, the difference in SNR with and without Galactic foreground decreases. This is because as $\rho^{\rm inj}$ increases, the turning point shifts to frequencies above $\sim 5\times 10^{-3}$Hz, where the Galactic foreground is suppressed. Consequently, the posteriors of $\rho$ with and without Galactic foreground overlap better as $\rho^{\rm inj}$ is increased, with virtually no difference in the case of $\rho^{\rm inj}\gtrsim 10^{-8}\mathrm{g}\,\mathrm{cm}^{-3}$. Additionally, the recovered SNR (with and without Galactic foreground) typically accumulates above $50$ only at $f\gtrsim(\mathrm{few})\times 10^{-3}$Hz. Thus, the measurements of signal parameters are typically informed by this specific frequency regime.
For each injection, we performed a separate set of runs with a vacuum SGWB model by setting $\alpha=\rho=0$ in $\Omega_{\rm RPLP}$. We compute the Bayes factor $\mathcal{B}^{\rm non-vac}_{\rm vac}$ in favor of the non-vacuum model ($\rho\neq 0$), given by Eq.˜27, which we list in Table˜2. For all injections, we find that $\mathcal{B}^{\rm non-vac}_{\rm vac}>10$, even when including the Galactic foreground. Thus, despite the marginalized posteriors for $A_{\rm vac}$ and $\rho$ displaying biases (compared to their vacuum asymptotic and injected values respectively), the RPLP model can detect dynamical friction effects by strongly disfavoring the vacuum model. We also checked that the posterior predictive from the RPLP model can accurately describe the injected signal within statistical errors, consistent with our Bayes factor results. We further discuss the biases (as seen in Fig.˜3) caused by the mismatch between the RPLP model and the true signal in Appendix C.
@@
If we detect a SGWB consistent with vacuum, our RPLP model can also be used to place upper bounds on the environmental parameters. In Fig.˜5, we show the constraints on $\rho$, with and without including the Galactic foreground parameters in our model. The constraints, as given by the $90\%$ one-sided credible interval, are informative, because the prior extends to $\rho=10^{-6}\mathrm{g\,cm^{-3}}$. When excluding the Galactic foreground, we find an upper bound of $\rho<8.51\times 10^{-11}\mathrm{g}\,\mathrm{cm}^{-3}$. Owing to the increase in dimensionality when including Galactic foreground, the constraint weakens slightly to $\rho<7.58\times 10^{-10}\mathrm{g}\,\mathrm{cm}^{-3}$.
Figure 6: Marginalized posterior probability for the vacuum model across different matter density regimes. Solid and dashed histograms represent the inference results with and without the inclusion of the Galactic foreground, respectively. Dash-dotted lines indicate the asymptotic value of $A_{\rm vac}$ (from Table˜1) with $\gamma=2/3$. Two-dimensional contours correspond to $90\%$ credible regions.
To explore systematic biases on the vacuum SGWB induced by neglecting environmental effects, we inject an SGWB containing dynamical friction effects, and analyze it with the vacuum model, i.e. $\Omega_{\rm RPL}$ with $\alpha=0$. In Fig.˜6, for the injected values of $\rho^{\rm inj}\in\{10^{-7},10^{-8},10^{-9},10^{-10},10^{-11}\}\mathrm{g}\,\mathrm{cm}^{-3}$, we show the recovered two- and one-dimensional marginalized posteriors for the vacuum amplitude $A_{\rm vac}$ and the vacuum spectral index $\gamma$. When assessing the systematic bias in $A_{\rm vac}$ and $\gamma$, we compare the maximum posterior point to the asymptotic vacuum values (vertical dash-dotted lines) listed in Table˜1. As expected, with increasing $\rho^{\rm inj}$, the systematic biases increase in significance. In more details, $\gamma$ is biased to values larger than the asymptotic value of $\gamma=2/3$, because the SGWB containing dynamical friction effects has a steeper slope, with an asymptotic value of 13/3. Due to the expected positive correlation with $A_{\rm vac}$ (given the functional form of the SGWB), we find that $A_{\rm vac}$ is biased to larger values. When inferring simultaneously on Galactic foreground parameters, due to the reduced constraining power on dynamical friction show in Fig.˜3, the biases on $A_{\rm vac}$ and $\gamma$ are milder, as expected.
@@ IV.2.2 Gas accretion
To characterize gas accretion measurability, we inject SGWB data (using models in Sec.˜III.2) with $f_{\rm Edd}^{\rm inj}\in\{0.01,0.10,1.00,10\}$. We use a log-uniform prior given by $\log_{10}f_{\rm Edd}\sim\mathcal{U}(-3,2)$, where the upper bound reflects astrophysical expectations of how large the Eddington ratio can be. We have checked that our results are robust to the exact choice of the lower bound, which we cannot set exactly to zero due to the log-uniform prior. We carry out the Bayesian inference just like we did for dynamical friction by fixing $\{\gamma,A_{m},\beta,\kappa\}$.
In Fig.˜7, we show the two-dimensional and one-dimensional marginalized posteriors of $A_{\rm vac}$ and $f_{\rm Edd}$. We accurately measure the vacuum amplitude across all injections, thus finding that marginalizing over gas accretion does not impact the measurability of the vacuum stellar SGWB. However, we do not obtain informative posteriors of Eddington ratios for any of the injected values, even with $f^{\rm inj}_{\rm Edd}=10$ marginally recovering the prior. The lack of constraining power on $f_{\rm Edd}$ is due to the turning point of the SGWB occurring at frequencies lower that the LISA sensitive band, as we anticipated qualitatively with BPLSs in Sec.˜III.2. Given our findings, we do not perform any additional Bayesian analysis for gas accretion effects.